Find the maximum horizontal range of a cricket ball projected with a velocity of 80m/s. If the ball is to have a range of 1003 meters, find the least angle of projection and the least time taken.

  1. Find the maximum horizontal range of a cricket ball projected with a velocity of 80m/s. If the ball is to have a range of 1003 meters, find the least angle of projection and the least time taken.

Grade:11

2 Answers

Unicorn
2 years ago
For maximum range angle of projection = 45
Range max=u^2/g=80^2/10=640 meters
 
If range =1003
1003=80^2 sin2 theta/g
Sin2 theta =1003/640
Theta=1/2 sin^-1(1003/640)
 
Time =2 u sintheta /g
::))
James M. Mason
one year ago
DO YOU PLAY LOTTERY?? 
 
I will assume you do or you wouldn't have ended up here on this webpage. But do you WIN the lottery? I will assume you don't, or you wouldn't have ended up here on this webpage. 
 
My name is James M. Mason, If you want a spell to win the lottery, look no further. Dr. Lucas lottery spell is amongst the most powerful spells I have experienced, and you will feel the power once He is done with your lottery spell. 
 
The spell uses a wide range of magical sources, then combines and funnels those energies into one single "beam" of luck attracting, destiny controlling, wealth manifesting pure magical power. 
 
Incorporating herbs, candles, the elements and much more simple magical techniques, this spell can be cast in a few minutes using a few very simple ingredients in His temple.
 
I am a living witness of his power because he just helped me with his lottery spell and I won CASH FOR LIFE LOTTERY, I'm so happy right now. Thanks to you Dr. Lucas. Contact Him through his private What's App Number +15302120104. 
Whichever type of lottery you play, HE can use His lottery spell to help you win.